# Pedagogical Alignment of Large Language Models

### v1

@@ 4 Experiments
In this section, we provide a comprehensive overview of our dataset construction, experimental design, and the subsequent findings derived from the application of state-of-the-art RLHF algorithms to train pedagogically-aligned LLMs.
@@ 4.1 Dataset and Evaluation
Our experimental design requires an extensive conversational dataset to train and test SFT and RLHF models effectively. The CLASS framework [9] used GPT-4 to generate 610 conversations to train their SFT model. However, we need to generate a larger volume of conversations to train RLHF models and conduct comprehensive testing. Following the strategy outlined by CLASS, we set about creating these additional conversations. We used the CLASS scaffolding prompts to generate the problems, subproblems, and associated hints, which were based on biology topics from the OpenStax college book [23], consistent with CLASS’ context. These problems seed the context of the conversations generated through GPT-4 in the next step using the CLASS conversational prompt. In these simulated student-tutor conversations, the tutor uses the subproblems and hints to guide the student toward the final answer. Through this methodology, we successfully generated an additional 1128 conversations, resulting in 1738 conversations. Each conversation, on average, comprises approximately 8.2 rounds and 1311.0 words. This comprehensive dataset provides a solid platform for training and comparing SFT and RLHF models.
@@
To create the training and testing set for SFT and RLHF, we randomly sampled our conversational dataset into 3 partitions: the SFT training dataset $D_{c}$ (600 conversations), the RLHF training dataset $D_{p}$ (600 conversations), and the test dataset $D_{t}$ (450 conversations). As each round within the conversations is processed into a separate data point, $D_{c}$ contains 4942 data points, $D_{p}$ contains 4921 data points, and $D_{t}$ contains 3701 data points.
@@
For the evaluation, we focus on the three classification fields within the model JSON responses described in section 3: Evaluation of Student Response, Action Based on Evaluation, and Subproblem State. We employ accuracy in these three fields to measure the pedagogical alignment of LLMs since these metrics offer a comprehensive view of the model’s performance, evaluating the correctness of the responses, the appropriateness of the actions, and the state of problem-solving. Hence, LLMs that achieve higher accuracy in these fields are considered to provide more pedagogically aligned student assistance.
@@
Table 2: SFT vs DPO, IPO, KTO. Each cell shows the average accuracy across three classification fields - Evaluation of student response, Action based on the evaluation, and Subproblem State, respectively. We observed that for both Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta), all three RL alignment algorithms - DPO, KTO, and IPO - showed notable accuracy gains in accuracy across all fields as compared to SFT. For both models, DPO and KTO outperformed IPO. IPO performed best for Vicuna (vicuna-7b-v1.5) compared to the other two.
Model SFT DPO IPO KTO
Mistral 35.0 (34,34,37) 74.0 (72,72,78) 57.7 (56,56,61) 74.7 (73,73,80)
Vicuna 74.3 (73,70,80) 75.0 (73,72,80) 76.0 (74,72,82) 75.3 (73,72,81)
Zephyr 10.7 (12,11,9) 74.0 (72,71,79) 21.7 (22,22,21) 74.7 (73,72,79)
Table 3: LLM Performance on key benchmarks. This table presents the performance of Mistral (Mistral-7B-Instruct-v0.2), Zephyr (zephyr-7b-beta), and Vicuna(vicuna-7b-v1.5) on different NLP benchmarks [24, 25]. We selected Mistral and Zephyr for our experiments based on their superior performance across various benchmarks, which provided a thorough and fair assessment across a spectrum of reasoning and general knowledge tasks. We chose Vicuna because it served as the base model for the CLASS framework.
Model Average ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K
Mistral 65.71 63.14 84.88 60.78 68.26 77.19 40.03
Zephyr 61.95 62.03 84.36 61.07 57.45 77.74 29.04
Vicuna 52.06 53.24 77.39 50.82 50.33 72.06 8.11
@@ 4.2 Models and and Training Details
We experimented on 3 different LLMs: Mistral (Mistral-7B-Instruct-v0.2), Zephyr (zephyr-7b-beta), and Vicuna (vicuna-7b-v1.5). We selected Mistral and Zephyr based on the their top performance on diverse NLP benchmarks (Table 3). Among all, several benchmark in particular evaluate scientific knowledge and education-based question answering capabilities: ARC (multiple choice questions from grade school science exams) [26], MMLU (multiple choice questions from 57 high school and college subjects across STEM, the humanities, the social sciences, and more) [27], and GSM8K (linguistically diverse grade school math word problems) [28]. These benchmarks offer a comprehensive, unbiased evaluation on the models’ performance, especially in education setting. Furthermore, despite Vicuna’s inferior performance compared to Mistral and Zephyr, we included it in our selection because it was the base model for the CLASS framework’s initial experiments and represented the state of the art at its launch.
@@
Our experiments included training and evaluating SFT and RLHF models. It took around 22 NVIDIA A6000 GPU hours for one cycle (SFT + Evaluation + RLHF + Evaluation). A total of 17 cycles, which are 374 GPU hours, are needed to complete all the experiments. We conducted the SFT experiments using FastChat [29]. For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna [29]: We employed a learning rate of 2e-5 with the AdamW optimizer [30], a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs. We conducted the RLHF experiments using TRL [31]. For the choice of RLHF hyperparamters, we referred to parameters used to align Zephyr [32]: We employed a learning rate of 1e-7 with the AdamW optimizer, a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
4.3 Experiment 1: SFT vs DPO, IPO, and KTO
@@
In this experiment, we compared three alignment algorithms: Direct Preference Optimization (DPO), Identity Preference Optimization (IPO), and Kahneman Tversky Optimization (KTO) on the SFT models. We measured the accuracy across three classification fields in the responses: Evaluation of Student Response, Action based on Evaluation, and Subproblem State.
@@
We used the same RLHF training setting as described in Section 4.2 with a beta of 0.1 for DPO, IPO, and KTO across all models. After applying the alignment algorithms, we observed significant accuracy improvements on Mistral and Zephyr SFT models. Specifically, Mistral demonstrated improvements of 39%, 22.7%, and 39.7% from DPO, IPO, and KTO respectively, while Zephyr showed improvements of 63.3%, 11%, and 64% respectively (Table 2). Notably, for both Mistral and Zephyr, DPO and KTO significantly outperformed IPO. However, the Vicuna model, already noted for its high SFT performance, showed only marginal improvements from the alignment; all three algorithms exhibited roughly similar performance. Our results are in alignment with the findings of similar studies [32], which have reported superior performance of DPO and KTO over IPO on Zephyr when the model was fine-tuned on preference datasets, such as Intel’s ORCA dataset pairs [33] and the Ultra-Feedback dataset [15].
@@
This experiment underscored the pivotal role of alignment algorithms in improving the accuracy of SFT models. Mistral and Zephyr, the top performing models, exhibited pronounced improvements over Vicuna. Therefore, we demonstrated the efficacy of pedagogical alignment on state of the art models.
4.4 Experiment 2: Effect of Beta on DPO and IPO
@@
Figure 3: DPO and IPO performance as a function of beta. This figure depicts how the beta hyper-parameter affects the average classification accuracy of DPO and IPO aligned models. We found that both DPO and IPO performances were significantly influenced by beta adjustments. Specifically, a beta of 0.3 optimized results for DPO, with Mistral and Zephyr achieving around 75% accuracy, while a higher beta of 0.9 notably reduced their accuracy. In contrast, IPO performed best with a beta of 0.1, and increasing the beta decreased the accuracy.
@@
Similar to experiments in Section 4.3, we measured the average accuracy across the three classification fields. We found that the performance of DPO and IPO is sensitive to the choice of beta. As depicted in figure 3, for DPO, we observed that a beta value of 0.3 yielded the best results: Mistral and Zephyr achieved an average accuracy of 74.7% and 75%. Conversely, a larger beta value, such as 0.9, diminished the performance of the models. Zephyr suffered a large loss of 25.7 % in average accuracy from using 0.9 instead of 0.3. As opposed to IPO, we observed that a beta of 0.1 produced the best performance (figure 3). A larger choice of beta in IPO almost always came with a loss of performance. As beta increases, Mistral’s accuracy dropped from 57.7% to 36 %, while Zephyr’s accuracy dropped from 21.7 % to 11.3 %. Our results corroborate the evidence from related studies [32], demonstrating that various alignment algorithms exhibit distinct sensitivities to beta values, necessitating tailored beta selections.
@@ 5 Conclusion
In this paper, we have introduced and explored the concept of pedagogical alignment of LLMs to align these models with educational objectives that are aimed at fostering optimal student learning. By reinterpreting the CLASS framework as a reinforcement learning problem, we constructed a reward dataset, thereby providing a foundation for enhancing the pedagogical alignment of LLMs. Our experimental results, derived from applying reinforcement learning-based alignment algorithms on state-of-the-art open-source LLMs, corroborate our hypothesis. The RLHF methods we employed have proven to outperform the SFT approach significantly, favoring the alignment of LLMs with pedagogical objectives. Furthermore, our study highlights the potential of online feedback for enhancing the performance of pedagogically-aligned LLMs. In conclusion, while our research has made significant strides in aligning LLMs with educational needs, further exploration and refinement of reinforcement learning methods for this purpose remain necessary. We anticipate that our work will serve as a stepping stone towards the optimal integration of LLMs in educational settings, paving the way for future research in this exciting domain.
Acknowledgments
This work was supported by NSF grants 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.

### v2

@@ 4 Experiments
In this section, we provide a comprehensive overview of our dataset construction, experimental design, and the subsequent findings derived from the application of state-of-the-art LHP algorithms to train pedagogically-aligned LLMs.
@@ 4.1 Dataset and Evaluation
Our experimental design requires an extensive conversational dataset to train and test SFT and LHP models effectively. The CLASS framework used GPT-4 to generate 610 conversations to train their SFT model. However, we need to generate a larger volume of conversations to train LHP models and conduct comprehensive testing. Following the strategy outlined by CLASS, we set about creating these additional conversations. We used the CLASS scaffolding prompts to generate the problems, subproblems, and associated hints, which were based on biology topics from the OpenStax college books [18], consistent with CLASS’ context. These problems seed the context of the conversations generated through GPT-4 in the next step using the CLASS conversational prompt. In these simulated student-tutor conversations, the tutor uses the subproblems and hints to guide the student toward the final answer. Through this methodology, we successfully generated an additional 1128 conversations, resulting in 1738 conversations. Each conversation, on average, comprises approximately 8.2 rounds and 1311.0 words. This comprehensive dataset provides a solid platform for training and comparing SFT and LHP models.
@@
To create the training and testing set for SFT and LHP, we randomly sampled our conversational dataset into 3 partitions: the SFT training dataset $D_{c}$ (600 conversations), the LHP training dataset $D_{p}$ (600 conversations), and the test dataset $D_{t}$ (450 conversations). As each round within the conversations is processed into a separate data point, $D_{c}$ contains 4942 data points, $D_{p}$ contains 4921 data points, and $D_{t}$ contains 3701 data points.
@@
For the evaluation, we focus on the three classification fields within the model JSON responses described in Section 3: Evaluation of Student Response, Action Based on Evaluation, and Subproblem State. We employ accuracy and F1 score in these three fields to measure the pedagogical alignment of LLMs since these metrics offer a comprehensive view of the model’s performance in evaluating the correctness of the responses (Evaluation of Student Response), the appropriateness of the actions (Action Based on Evaluation), and the state of problem-solving (Subproblem State). Hence, LLMs that achieve higher accuracy and F1 score in these fields are considered to provide more pedagogically aligned assitance to students.
Model Metric SFT DPO IPO KTO
Mistral Acc 35.0 (34,34,37) 74.0 (72,72,78) 57.7 (56,56,61) 75.3 (73,73,80)
F1 0.34 (0.42,0.24,0.37) 0.56 (0.63,0.37,0.67) 0.48 (0.56,0.33,0.56) 0.55 (0.62,0.37,0.65)
Zephyr Acc 10.7 (12,11,9) 74.0 (72,71,79) 21.7 (22,22,21) 74.7 (73,72,79)
F1 0.14 (0.21,0.12,0.10) 0.56 (0.65,0.38,0.66) 0.26 (0.32,0.19,0.28) 0.54 (0.63,0.37,0.63)
@@
Table 2: SFT vs DPO, IPO, KTO. Each cell shows the average accuracy and average F1 score across three classification fields - Evaluation of student response, Action based on the evaluation, and Subproblem State, respectively. We observed that for both Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta), all three alignment algorithms - DPO, KTO, and IPO - showed notable gains in accuracy and F1 score across all fields as compared to SFT. Across both models, DPO and KTO exhibited comparable performance, outperforming IPO.
@@ 4.2 Models and and Training Details
We experimented on two different LLMs: Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta). We selected Mistral and Zephyr based on their top performance on diverse NLP benchmarks on OpenLLM leaderboard [19, 20] which evaluate scientific knowledge and education-based question answering capabilities: ARC (reasoning based science questions) [21], MMLU (multiple choice questions from 57 high school and college subjects across STEM, the humanities, the social sciences, and more) [22], and GSM8K (grade school math word problems) [23]. These benchmarks offer a comprehensive evaluation on the models’ performance, especially in education setting.
@@
Our experiments included training and evaluating SFT and LHP models. It took around 22 NVIDIA A6000 GPU hours for one cycle (SFT + Evaluation + LHP + Evaluation). A total of 17 cycles, which are 374 GPU hours, are needed to complete all the experiments. We conducted the SFT experiments using FastChat [24]. For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna [24]: We employed a learning rate of 2e-5 with the AdamW optimizer [25], a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs. We conducted the LHP experiments using TRL [26]. For the choice of LHP hyperparamters, we referred to parameters used to align Zephyr [27]: We employed a learning rate of 1e-7 with the AdamW optimizer, a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
4.3 Main findings: SFT vs LHP
@@
In this experiment, we compared three alignment algorithms: Direct Preference Optimization (DPO), Identity Preference Optimization (IPO), and Kahneman Tversky Optimization (KTO) on the SFT models. We measured the accuracy and F1 score across three classification fields in the responses: Evaluation of Student Response, Action based on Evaluation, and Subproblem State.
@@
We used the same LHP training setting as described in Section 4.2 with a beta of 0.1 for DPO, IPO, and KTO across all models. After applying the alignment algorithms, we observed significant performance improvements on Mistral and Zephyr SFT models. Specifically, Mistral demonstrated accuracy increase of 39%, 22.7%, and 39.7% from DPO, IPO, and KTO respectively, while Zephyr showed improvements of 63.3%, 11%, and 64% respectively (Table 2). We observed similar increase in F1 score as well: Mistral showed increase of 0.22, 0.14, and 0.21 from DPO, IPO, and KTO, and Zephyr showed improvements of 0.42, 0.12, and 0.4 respectively. Notably, for both Mistral and Zephyr, DPO and KTO significantly outperformed IPO.
Our results are in alignment with the findings of similar studies [27], which have reported superior performance of DPO and KTO over IPO on Zephyr when the model was fine-tuned on preference datasets, such as Intel’s ORCA dataset pairs [28] and the Ultra-Feedback dataset [14].
@@
This experiment underscored the pivotal role of alignment algorithms in improving the accuracy and F1 score of SFT models. Mistral and Zephyr, the top performing models, exhibited pronounced improvements using LHP algorithms compared to SFT.
4.4 Examining LHP through lens of perplexity
For understanding the effectiveness of LHP algorithms compared to SFT, we propose a second preference dataset generation technique. This technique specifically allows us to perform a perplexity analysis to compare the LLM’s tendency to provide hints and guidance versus direct responses.
The essence of pedagogical alignment is the LLM’s propensity to offer guidance (as seen in Actions 1 and 4) over direct solutions (as in Actions 2 and 5). By this definition, Action 2 can never precede Action 1, and likewise, Action 5 can never precede Action 4 (for more details about actions please refer to Table 1). With this critical insight, we use GPT-4 to generate pedagogically misaligned Actions (Action 2 for the first occurrence of Action 1 and Action 5 for the first occurrence of Action 4). This approach allows us to compare the perplexity of responses corresponding to Action 1 and its misaligned counterpart (Action 2), and similarly for Actions 4 and 5.
Building upon the test split (450 conversations) of the originally generated dataset, we generate an alternative opposite action (Action 2 for Action 1, and Action 5 for Action 4) for the first occurrence of Action 1 or 4 in each conversation. We then use this newly constructed Action dataset to measure the perplexity of responses from the original instruct model (Base), its SFT fine-tuned counterpart (SFT), and its LHP version (DPO) (Table 3).
Our analysis reveals that for both Mistral and Zephyr Base models, responses with Action 1 have higher perplexity than those with Action 2, and likewise for Actions 4 and 5. This suggests that the base models are more inclined to provide direct solutions than offer hints and guidance. The SFT process adjusts this behavior by increasing the perplexity of responses with Actions 2 and 5, thus making the models more likely to produce responses with Actions 1 and 4. However, SFT alone doesn’t fully optimize this behavior. As seen in the perplexity results for Mistral and Zephyr SFT models (Table 3), there is only a minimal decrease or even a slight increase in the perplexity of Action 1 and 4. This indicates that SFT, while useful, doesn’t significantly shift the model’s behavior towards providing more guidance over direct solutions.
In contrast, LHP dramatically optimizes this behavior. For the Mistral model, the perplexity of Action 1 and 4 decreases from 2.05 and 2.14 to 1.89 and 1.96 (SFT vs DPO) respectively. Similarly, for the Zephyr model, the perplexity of Action 1 and 4 decreases from 2.09 and 2.18 (SFT) to 1.97 and 2.05 (DPO) respectively. This decrease in perplexity indicates a higher likelihood of the model providing guidance over direct solutions.
Simultaneously, the perplexity of Action 2 and 5 increases and from 13.9 and 13.06 (Zephyr SFT) to 14.9 and 14.38 (Zephyr DPO). This increase in perplexity indicates a lower likelihood of the model providing direct solutions, further emphasizing the model’s shift towards providing guidance.
Model A1 A2 A4 A5
Mistral (Base) 2.11 2.06 2.00 1.97
Mistral (SFT) 2.05 5.55 2.14 5.40
Mistral (DPO) 1.89 5.48 1.96 5.34
Zephyr (Base) 2.05 2.09 2.03 2.07
Zephyr (SFT) 2.09 13.9 2.18 13.06
Zephyr (DPO) 1.97 14.9 2.05 14.38
Table 3: Perplexity of responses. Perplexity analysis reveals that the base models of Mistral and Zephyr are more inclined to provide direct solutions (Action 2 and 5) than guidance and hints (Action 1 and 4). The SFT process partially shifts this behavior, increasing the perplexity on responses with Action 2 and 5, thus making the models more likely to produce responses with Action 1 and 4. However, the SFT process alone doesn’t fully optimize this behavior, as seen by the minimal decrease or slight increase in the perplexity of Action 1 and 4. In contrast, DPO dramatically optimizes this behavior by reducing perplexity on Action 1 and 4 while simultaneously increasing perplexity on Action 2 and 5. This shift in perplexity indicates a significant behavioral shift in the LLM, making it more aligned with the role of a tutorbot by favoring guidance over direct solutions. For more details about actions please refer to Table 1.
4.5 Effect of Beta on DPO and IPO
@@
Figure 3: DPO and IPO performance as a function of beta. We found that classification accuracy and F1 score using both DPO and IPO were significantly influenced by beta hyper-parameter adjustments. Specifically, a beta of 0.3 optimized results for DPO, with Mistral and Zephyr achieving around 75% accuracy and 0.56 F1 score, while a higher beta of 0.9 notably reduced accuracy and F1 score. In contrast, IPO performed best with a beta of 0.1 and increasing beta decreased accuracy and F1 score.
@@
Similar to experiments in Section 4.3, we measured the average accuracy and F1 score across the three classification fields. We found that the performance of DPO and IPO is sensitive to the choice of beta. As depicted in Figure 3, for DPO, we observed that a beta value of 0.3 yielded the best results: Mistral and Zephyr achieved an average accuracy of 74.7% and 75% and an average F1 score of 0.55 and 0.56. Conversely, a larger beta value, such as 0.9, diminished the performance of the models. Zephyr suffered a large loss of 25.7 % in average accuracy from using 0.9 instead of 0.3, and a loss of 0.13 in average F1 score. As opposed to IPO, we observed that a beta of 0.1 produced the best performance (Figure 3). A larger choice of beta in IPO almost always came with a loss of performance. As beta increased, Mistral’s accuracy dropped from 57.7% to 36 %, and its F1 score decreased from 0.55 to 0.49, while Zephyr’s accuracy dropped from 21.7 % to 11.3 %, and its F1 score decreased from 0.56 to 0.43. Our results corroborate the evidence from related studies [27], demonstrating that various alignment algorithms exhibit distinct sensitivities to beta values, necessitating tailored beta selections.
@@ 5 Conclusion
In this paper, we have investigated the application of learning from human preferences (LHP) methods to align LLMs with educational objectives, aiming to foster optimal student learning. By constructing a preference dataset specifically designed for pedagogical alignment, we have laid the foundation for enhancing the effectiveness of LLMs in educational contexts. Our experimental results, derived from applying LHP alignment algorithms on state-of-the-art open-source LLMs, demonstrate the superiority of LHP methods over the traditional supervised fine-tuning (SFT) approach, significantly improving the alignment of LLMs with pedagogical objectives. Our another key contribution is the development of a novel approach for evaluating the pedagogical effectiveness of LLMs. By curating a second preference dataset that compares the perplexity of responses offering scaffolding versus those providing direct solutions, we have introduced a new methodology that quantifies the extent to which LLMs prioritize pedagogically effective actions. This approach opens up new possibilities for assessing and optimizing the performance of LLMs in educational settings, providing a valuable tool for researchers and practitioners alike.